OpenStudy (anonymous):

1.Solve the system of equations using linear combination.

OpenStudy (anonymous):

9+c=9 8a+4.5c=58

OpenStudy (anonymous):

simultaneous equations?

OpenStudy (anonymous):

oh gosh i have no idea....

OpenStudy (anonymous):

yea they r the same ok ill show u

OpenStudy (anonymous):

ok ok thanks! :)

OpenStudy (anonymous):

9+c=9 c=0 therefore 8a+4.5(0)=58 8a=58 a=7.25

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

yea ok i see now thank you :)

OpenStudy (anonymous):

:)